

********************************************************************
UARS WEEKLY STATUS REPORT - 15 NOV 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
Despite the unexpected UARS shutdown, SUSIM successfully gathered all
of its daily mid-resolution and MgII solar scans this past week
during periods of predicted UARS telemetry.
***********
ACRIM
***********
The ACRIM2 Science Team has submitted FOT control sequences designed to
test the ability to stimulate instrument turn-on by switching power
busses A and B off and on. Reproduction of the power-on that occured after
the 8/01 TMON is sought.
***********
PEM
***********
The PEM web site has moved. It has a new URL. Access the PEM web site
using http://wwwpem.space.swri.edu. The PEM data from the GUVI conjunction
periods has moved to http://wwwpem.space.swri.edu/guvi-pem.html.
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, all HEPSA data between 1991274 and 1992119 has been
reprocessed and is archived in the DAAC. HEPSA data for 1992120, between
1992122 and 1992128, and between 1992130 and 1992199 have been reprocessed
and waiting DAAC archival.
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes when both instruments saw aurora.
PEM has seen no aurora during the second conjunction period. PEM data is
being processed for the third conjunction period.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG assisted with troubleshooting the problem
of DECNET communications between instrument Remote
Analysis Computers and the UARS Command Management
System until the problem was resolved on 13 November.
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 02 November 2002 (Orbit 60940, GMT Day 306) through
08 November 2002 (Orbit 61044, GMT Day 312).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 308/1941 60982 POWER ON
PEM 309/2309 60999 AXIS POWER OFF
HALOE 310/1254 61008 POWER ON
The beta angle ranged between the angles of 65.5 degrees to 44.6
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 19 November 2002 (DOY 323).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.28 degrees C. Additional
battery performance data for each battery on 01 November and 08
November 2002 are provided in Appendix D.
The clock error ranged from -12.0 to 10.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
HRDI = 02
PEM = 01
HALOE = 02
WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Two (2) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT is happy to report that we have successfully recovered from
the OBC halt anomaly which occurred on 21 October 2002. All
instruments that are scheduled on are now back on including HRDI. The
FOT observed a HRDI HPC protection telemetry point toggling at a
fairly rapid pace before the anomaly. The HRDI PI team was notified
of this and investigated this circuitry. They found that this
signature occurs nominally and did not believe this to have caused the
OBC halt. Data was pulled by the FOT which confirmed that signature
is nominal. HRDI was powered back on on the 308/1941.
A full analysis of the OBC hardware dumps, which were taken
immediately after the anomaly, has not revealed any indications of any
abnormal behavior. The FOT is planning to try to load this image to
the simulator in order to discover more about the state of the OBC
during the anomaly. If this test does not show any clues about the
problem then the cause of the anomaly may remain a question mark (or
be attributed to a single event upset). The investigation continues
but as of right now there are no firm leads to follow.
MEETINGS
The FOT attended the UARS Monthly Status Review on 06 November 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 November 2002 (DOY 322).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 November 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60973 308/0603 02-424 OBC/CLK CLK ADJ 20.6 TO 0.00 MSEC/DAY
60982 308/1941 02-425 HRDI POWER ON
60986 309/0200 02-426 HRDI PB TELE ON
60994 309/1439 02-427 TAM TAM REV ALIGN
60999 309/2309 02-428 PEM AXIS POWER OFF
61003 310/0600 02-429 OBC/CLK CLK ADJ 0.00 TO 20.6 MSEC/DAY
61008 310/1254 02-430 HALOE POWER ON
61028 311/2121 02-431 OBC/CLK CLK ADJ 20.6 TO 0.00 MSEC/DAY
61039 312/1532 02-432 WINDII WINDII ERR CLEAR
61042 312/1932 02-433 HALOE DAP CMDS
61043 312/2045 02-434 OBC SOFT BANK DUMPS
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 November 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 November 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60973 308/0605 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
61003 310/0602 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
61028 311/2131 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 November 2002
01 November 2002 - GMT Day 304 Beta = 68.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -145.6/-179.2 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +05.6/-33.6 +3.23/+1.97 +19.2/-16.0 26.6 1.040 15.5
3 +16.8/-22.4 +2.60/+1.34 +18.4/-15.2 26.4 1.030 15.0
08 November 2002 - GMT Day 312 Beta = 44.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -151.2/-179.2 -3.05/-4.62 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-44.8 +5.47/+4.51 +27.2/-15.2 24.8 1.040 20.5
3 +22.4/-22.4 +4.51/+3.23 +26.4/-15.2 24.8 1.030 20.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
- None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 November 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-053 FOT 60630 285/1008 30382 GENERIC LATE ACQUISITION
02-056 FOT 60775 295/0122 OBC CRASH